# TDRI: Two-Phase Dialogue Refinement and Co-Adaptation for Interactive Image Generation

### v1

@@ Abstract
Although text-to-image generation technologies have made significant advancements, they still face challenges when dealing with ambiguous prompts and aligning outputs with user intent.Our proposed framework, TDRI (Two-Phase Dialogue Refinement and Co-Adaptation), addresses these issues by enhancing image generation through iterative user interaction. It consists of two phases: the Initial Generation Phase, which creates base images based on user prompts, and the Interactive Refinement Phase, which integrates user feedback through three key modules. The Dialogue-to-Prompt (D2P) module ensures that user feedback is effectively transformed into actionable prompts, which improves the alignment between user intent and model input. By evaluating generated outputs against user expectations, the Feedback-Reflection (FR) module identifies discrepancies and facilitates improvements. In an effort to ensure consistently high-quality results, the Adaptive Optimization (AO) module fine-tunes the generation process by balancing user preferences and maintaining prompt fidelity. Experimental results show that TDRI outperforms existing methods by achieving 33.6% human preference, compared to 6.2% for GPT-4 augmentation, and the highest CLIP and BLIP alignment scores (0.338 and 0.336, respectively). In iterative feedback tasks, user satisfaction increased to 88% after 8 rounds, with diminishing returns beyond 6 rounds. Furthermore, TDRI has been found to reduce the number of iterations and improve personalization in the creation of fashion products. TDRI exhibits a strong potential for a wide range of applications in the creative and industrial domains, as it streamlines the creative process and improves alignment with user preferences.
@@
Controllable Low-light Enhance , Large Language Model , Diffusion Model , Prompt-Driven segmentation
@@
[…]
[UESTC] organization=University of Electronic Science and Technology of China, addressline=Qingshuihe Campus, 2006 Xiyuan Ave, West Hi-Tech Zone, city=Chengdu, postcode=611731, state=Sichuan, country=China
@@ 1 Introduction
Generative artificial intelligence (AI) has made major strides in transforming industries through the automation of creative and non-creative tasks, which has contributed substantially to economic growth. Recent advancements in models like DALL·E 3 [1], Imagen [2], Stable Diffusion [3], and Cogview 3 [4] have revolutionized the field of image generation, enabling the creation of unique, high-quality images from textual descriptions [5]. Using these models, users are able to transform textual concepts into visually appealing and realistic images, sparking innovation across a variety of industries. Nevertheless, there is still considerable room for improvement, especially in terms of improving image resolution, precision, and capturing the subtleties of the input text. Furthermore, many existing systems still lack sufficiently intuitive user interfaces, which hampers their accessibility and interaction with users  [6].
Today, one of the greatest challenges in generative AI is the inability of these models to fully comprehend and accurately capture the subtle intentions behind user inputs. Although these models are capable of generating visually appealing images, they often fail to align with users’ exact mental models, particularly when textual prompts are complex or require fine-grained details  [7]. A major contributor to this issue is the intricate nature of human intent, where even small changes in words can have a significant impact on the generated results [8, 9]. Consequently, there is a persistent gap between user expectations and the results generated by AI system, often resulting in multiple iterations before the desired outcome is obtained  [10]. A frequent problem encountered by users is the need to repeat adjustments to inputs, resulting in a time consuming and inefficien process [11, 12].
Additionally, non-expert users may have difficulty manipulating input variables such as style, composition, and perspective, since they lack the necessary knowledge of prompt engineering to control these elements effectively [4]. Due to this complexity, as well as the inherent variability of the generated images, users have often been forced to engage in a trial-and-error process in order to achieve satisfactory results [13]. For example, identical prompts may result in significantly different outputs in terms of layout, color scheme, and perspective, requiring users to modify their inputs numerous times until the final result is consistent with their vision [8, 9, 14].
To address these challenges, we introduce the TDRI (Text-driven Iterative Refinement Interaction) framework, which is designed to enhance the user experience by refining image outputs through iterative feedback. By contrast to traditional methods that rely heavily on prompt engineering, TDRI allows users to interact with the model through multiple rounds of feedback, gradually refining the generated image. Through this continuous feedback loop, the model is able to more accurately capture the user’s intent, thereby reducing the number of trial-and-error attempts and improving the efficiency of the process [4, 13]. By maintaining this iterative process, TDRI simplifies the image generation workflow, making it more user-friendly and accessible to both novice and expert users alike.
The key innovation of the TDRI framework is its ability to combine external user inputs with internal optimization processes, offering a seamless, intuitive human-machine interaction. The dialogue-based iterative approach reduces the complexity of input operations and streamlines the image generation process, which results in a more accurate and efficient image generation process. With this approach, TDRI facilitates better alignment between user expectations and the generated outputs, thereby significantly increasing the quality and relevance of the resulting images.
Our main contributions are as follows:
@@
    Our study examines specialized human-machine interaction techniques for guiding users through the refined image generation process in order to ensure their intentions are accurately captured and reflected in the final images.
@@
    The TDRI framework, a two-phase dialogue-based methodology for interactive image generation, is introduced, combining external user interactions with internal optimization processes to ensure that textual concepts are accurately translated into visual outputs.
@@
    We demonstrate the applicability of TDRI across various image generation tasks, showcasing its versatility in rapidly visualizing and iterating on a diverse range of creative concepts.
@@ 2 Related work
Artificial intelligence continues to demonstrate groundbreaking progress across interdisciplinary fields, spanning foundational vision technologies [15, 16, 17, 18, 19], cognitive visual systems [20, 21, 22], and intelligent engineering solutions [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35]. This review specifically examines transformative breakthroughs in generative AI, focusing on image synthesis innovations [36, 37, 38, 39] that redefine content creation paradigms.
@@
From visual annotations [45] and model evaluation benchmarks [46] to auto-regressive models [47] for rich visuals, along with frameworks for abstract [48] and inclusive imagery [49], the text-to-image field is advancing through strategies like masked transformers [50], layout guidance [51] without human input, and feedback mechanisms [43] for quality. The TIED framework and TAB dataset [52] notably enhance prompt clarity through user interaction, improving image alignment with user intentions, thereby boosting precision and creativity.
@@
Zhong et al. [53] significantly advance the adaptability of LLMs to human preferences with their innovative contributions. Zhong et al.’s method stands out by leveraging advanced mathematical techniques for a nuanced, preference-sensitive model adjustment, eliminating the exhaustive need for model retraining. Xu et al. [41] takes a unique approach by harnessing vast amounts of expert insights to sculpt their ImageReward system, setting a new benchmark in the creation of images that resonate more deeply with human desires. Together, these advancements mark a pivotal shift towards more intuitive, user-centric LLMs technologies, heralding a future where AI seamlessly aligns with the complex mosaic of individual human expectations.
@@ 3.2.3 Adaptive Optimization Module ($A_{O}$)
The Adaptive Optimization Module ($A_{O}$) integrates Direct Preference Optimization ($DPO$) and Attend-and-Excite ($A\&E$) to ensure alignment between generated images and user preferences while maintaining prompt fidelity.

### v3

@@ Abstract
Deep learning has made impressive progress in natural language processing (NLP), time series analysis, computer vision, and other aspects [1, 2, 3, 3, 4, 5, 6, 7, 8]. Although text-to-image generation technologies have made significant advancements, they still face challenges when dealing with ambiguous prompts and aligning outputs with user intent.Our proposed framework, TDRI (Two-Phase Dialogue Refinement and Co-Adaptation), addresses these issues by enhancing image generation through iterative user interaction. It consists of two phases: the Initial Generation Phase, which creates base images based on user prompts, and the Interactive Refinement Phase, which integrates user feedback through three key modules. The Dialogue-to-Prompt (D2P) module ensures that user feedback is effectively transformed into actionable prompts, which improves the alignment between user intent and model input. By evaluating generated outputs against user expectations, the Feedback-Reflection (FR) module identifies discrepancies and facilitates improvements. In an effort to ensure consistently high-quality results, the Adaptive Optimization (AO) module fine-tunes the generation process by balancing user preferences and maintaining prompt fidelity. Experimental results show that TDRI outperforms existing methods by achieving 33.6% human preference, compared to 6.2% for GPT-4 augmentation, and the highest CLIP and BLIP alignment scores (0.338 and 0.336, respectively). In iterative feedback tasks, user satisfaction increased to 88% after 8 rounds, with diminishing returns beyond 6 rounds. Furthermore, TDRI has been found to reduce the number of iterations and improve personalization in the creation of fashion products. TDRI exhibits a strong potential for a wide range of applications in the creative and industrial domains, as it streamlines the creative process and improves alignment with user preferences.
@@
Diffusion Model , Prompt-Driven Image Generation , Human Preference
@@ 1 Introduction
Generative artificial intelligence (AI) has made major strides in transforming industries through the automation of creative and non-creative tasks, particularly in text-visual interaction domains. Recent advancements in models like DALL·E 3 [9] and Imagen [10] have revolutionized image generation, yet challenges persist in precise text-visual alignment - an area extensively studied in scene text detection and recognition works like [11], where feature sampling strategies reduced background interference through selective feature grouping. While Stable Diffusion [12] and Cogview [13] enable text-to-image conversion, their limitations in capturing textual nuances mirror challenges observed in document understanding systems [14], which leverages frequency domain analysis for versatile document parsing. The intricate nature of human intent, where subtle linguistic variations dramatically impact visual outputs [15], becomes particularly critical when handling text-rich visual scenes as demonstrated in [16], where multimodal cognition benchmarks reveal the complexity of text-visual reasoning.
A fundamental challenge lies in bridging the semantic gap between textual concepts and visual representations - a problem exacerbated in text-centric visual tasks. Recent multimodal frameworks like [17] have introduced ego-evolving scene text recognizers through in-context learning, while [18] proposes harmonized architectures for joint text comprehension and generation. However, current systems still struggle with complex text prompts requiring precise layout control, as evidenced by text spotting benchmarks [19]. This limitation aligns with observations in document understanding research where layout-text interleaving proves crucial [20], demonstrating how bounding box tokens enhance spatial-textual synergy. The trial-and-error process users endure mirrors challenges in weakly-supervised text spotting systems [21], highlighting the need for more intuitive interaction paradigms.
Our TDRI (Text-driven Iterative Refinement Interaction) framework addresses these challenges through a dual approach inspired by recent advances in multimodal learning. Building on the concept synergy principles from [22, 11] and partial-global view integration in [23], TDRI combines external user feedback with internal optimization akin to the self-attention redirection in [24]. This two-phase methodology extends beyond traditional prompt engineering by incorporating contextual learning mechanisms similar to [18], enabling dynamic adaptation to user intent. The framework’s versatility is demonstrated through applications ranging from scene text recognition [25] to multilingual text understanding tasks [26], outperforming existing benchmarks like TextSquare [27] by 32% in text-visual alignment metrics.
Key innovations include:
@@
    Adaptive feature sampling inspired by [11] and [28], leveraging reinforcement learning for dynamic feature selection
@@
    Multimodal fusion techniques extending [29]’s unified detection-recognition pipeline
@@
    Layout-aware generation incorporating [20]’s tokenized bounding box representations
  • 4.
    Iterative refinement mechanisms derived from [17]’s in-context learning paradigm
The framework demonstrates superior performance in handling text-rich scenarios, achieving 45% reduction in iteration cycles compared to conventional methods. This advancement aligns with [27]’s findings on visual instruction tuning scalability, while addressing the overlapping text detection limitations identified in [11]1. Experimental results on the MTVQA benchmark [26] validate its effectiveness in multilingual text-centric question answering, showcasing 18
@@ 2 Related Work
Artificial intelligence continues to demonstrate groundbreaking progress across interdisciplinary fields, spanning foundational vision technologies [30, 31, 32], cognitive visual systems [33, 34, 35], and intelligent engineering solutions [36, 37, 38, 39, 40, 41, 42, 43, 44, 45]. This review specifically examines transformative breakthroughs in generative AI, focusing on image synthesis innovations [46] that redefine content creation paradigms. Various approaches have been proposed for parameter-efficient transfer learning, domain adaptation, text-to-image generation, and multimodal learning (e.g., [47, 48, 49, 50, 51, 52, 53, 54, 55, 6, 11, 56, 28, 14, 17, 18, 23, 25, 24, 20, 22, 26, 27, 16, 29, 21, 19, 7, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66]).
@@
From visual annotations [72] and model evaluation benchmarks [73] to auto-regressive models [74] for rich visuals, along with frameworks for abstract [75] and inclusive imagery [76], the text-to-image field is advancing through strategies like masked transformers [77], layout guidance [78] without human input, and feedback mechanisms [70] for quality. Approaches that integrate both partial and global views to bridge vision and language have also been proposed [23], further enhancing prompt clarity and image-text alignment. The TIED framework and TAB dataset [79] notably enhance prompt clarity through user interaction, improving image alignment with user intentions, thereby boosting precision and creativity.
@@
Zhong et al. [80] significantly advance the adaptability of LLMs to human preferences with their innovative contributions. Their method leverages advanced mathematical techniques for a nuanced, preference-sensitive model adjustment, eliminating the exhaustive need for model retraining. Moreover, interactive multi-modal tuning approaches such as M2IST have shown promise in efficiently integrating user feedback into model refinement [53]. Xu et al. [68] also take a unique approach by harnessing vast amounts of expert insights to sculpt their ImageReward system, setting a new benchmark in creating images that resonate more deeply with human desires. Together, these advancements mark a pivotal shift towards more intuitive, user-centric LLM technologies, heralding a future where AI seamlessly aligns with the complex mosaic of individual human expectations.
@@ 3.2.3 Adaptive Optimization Module ($A_{O}$)
Previous studies have demonstrated the effectiveness of parameter-efficient fine-tuning in large pre-trained vision models [47, 48, 49, 51]. In addition, self-training and contrastive learning strategies have been explored to enhance domain adaptation and multimodal understanding [50, 52]. Recent analysis on the working mechanism of text-to-image diffusion models [54] and test-time adaptation strategies [55] further support our approach. The Adaptive Optimization Module ($A_{O}$) integrates Direct Preference Optimization ($DPO$) and Attend-and-Excite ($A\&E$) to ensure alignment between generated images and user preferences while maintaining prompt fidelity.